OpenStudy (anonymous):

Evaluate 9P5. 15,120 45 362,880

OpenStudy (anonymous):

In general\[P(n,k) = \frac{n!}{(n-k)!}\]Here you have n=9, k=5

OpenStudy (anonymous):

45?

OpenStudy (anonymous):

It would be\[\frac{9!}{(9-5)!} = \frac{9!}{4!} = \frac{1*2*3*4*5*6*7*8*9}{1*2*3*4}\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Did you find the answer?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

no?

OpenStudy (anonymous):

i am really confused

OpenStudy (anonymous):

Simplify this \[\frac{1*2*3*4*5*6*7*8*9}{1*2*3*4}\] where * denotes multiplication. This can't get any simpler.

OpenStudy (anonymous):

k thanks i got it now

OpenStudy (anonymous):

its the last one

OpenStudy (anonymous):

If you do the multiplication, you will find the correct answer.